Rate of formation of So3 according to the reaction 2So2 + O2 gives 2So3 is 1.6×10^-3 kg/min hence rate at which So2 reacts as.
1
Expert's answer
2019-08-28T04:27:00-0400

"2 SO_2+O_2 \to 2 SO_3"

By rate law kinetics:

"-\frac{1}{2} \frac{d[SO_2]}{dt}=\frac{1}{2} \frac{d[SO_3]}{dt}"

hence,"- \frac{d[SO_2]}{dt}" "=" "\frac{d[SO_3]}{dt}"

so,rate at which "SO_2" reacts="-(" rate at which "SO_3" forms.")"

"=-1.6 \times 10^{-3} \space kg /min" (negative sign represents the diminishing concentration of "SO_2" )